OpenStudy (anonymous):

An object moves along the x-axis, its position at each time t> 0 given by X(t)= 1/4t^4- t^3 + t^2 Determine the time interval(s), if any, during which the object moves right.

OpenStudy (anonymous):

When \(x'(t) > 0\) it is moving right.

OpenStudy (anonymous):

Do you know how to find derivatives @kizi239 ?

OpenStudy (anonymous):

but i need to find the time intervals

OpenStudy (anonymous):

Do you know how to find derivatives, yes or no? Don't worry, I'll walk you through it, but I need to know what you can do.

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

To give you a bit more confidence, remember that the solution to an inequality is an interval.

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

For example, the inequalty \(x < 2\) is the interval \((-\infty, 2)\)

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

So first of all, what is \(x'(t)\)?

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

Well, we are looking for when \(x'(t) = 0\), what I mean is, what is the derivative of \(x(t)\). Remember that \(x(t) =\frac{1}{4}t^4- t^3 + t^2\)

OpenStudy (anonymous):

@kizi239 So can you take the derivative?

OpenStudy (anonymous):

[t ^{3}-3t ^{2} + 2t\]

OpenStudy (anonymous):

Great. Now when we want to factor this function. If we can get it into the form: \[ (t-r_1)(t-r_2)(t-r_3) < 0 \]It'll be easier to find the intervals.

OpenStudy (anonymous):

\[ x'(t) = t ^{3}-3t ^{2} + 2t < 0 \\ (t^{2}-3t + 2)(t) < 0 \]We want to use the quadratic formula to find the roots of the polynomial \(t^{2}-3t + 2\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@kizi239 Will you do it?

OpenStudy (anonymous):

-2 and 8

OpenStudy (anonymous):

Wait, does that make sense? \(-2 + 8 = 6 \quad -2 \times 8 = -16\). Ummmm

OpenStudy (anonymous):

\[ \frac{-(-3)\pm\sqrt{(-3)^2-4(1)(2)}}{2(1)} \]

OpenStudy (anonymous):

thats wat i did

OpenStudy (anonymous):

\[ \begin{split} \frac{-(-3)\pm\sqrt{(-3)^2-4(1)(2)}}{2(1)} &=\frac{3\pm\sqrt{9-8}}{2} \\ &=\frac{3\pm1}{2} \\ &=\frac{4}{2},\quad \frac{2}{2} \\ &= 2, \quad 1 \end{split} \]

OpenStudy (anonymous):

Maybe you messed up somehow.

OpenStudy (anonymous):

ohh i made a mistake srry

OpenStudy (anonymous):

So anyway, we end up with: \[ (t-2)(t-1)(t) <0 \]

OpenStudy (anonymous):

txs so much

OpenStudy (anonymous):

Forgive me, I made a mistake... it should be \[ (t-2)(t-2)(t) > 0 \]

OpenStudy (anonymous):

ok txs

OpenStudy (anonymous):

that 2 should be a 1, sorry somehow I'm messing up. we're still not done yet though!

OpenStudy (anonymous):

So we are multiplying three factors: \((t-1)\), \((t-2)\), and \(t\). We want them to be greater than \(0\). So we have to remember our rules about multiplying negatives positives... We want either all of the factors to be positive. Or we want one factor to be positive and two to be negative. The two negatives will cancel out.

OpenStudy (anonymous):

They told us at the beginning that \(t>0\), so we know that the \(t\) factor will ALWAYS be positive.

OpenStudy (anonymous):

So we want the remaining factors to both be positive or both be negative. Basically: \[ t-1 > 0, \quad t-2 > 0 \]O: \[ t-1 < 0, \quad t-2 < 0 \]Does that make sense?

OpenStudy (anonymous):

This gives us \[ t > 1, \quad t > 2 \]Which is just \(t > 2\). Or \[ t < 1, \quad t < 2 \]Which is just \(t < 1\).

OpenStudy (anonymous):

That gives us 4 inequalities all together: \[ t > 0 \\ t < 1 \\ t > 2 \\ t < \infty \]Which is just \[ 0 < t < 1 \\ 2 < t < \infty \]In interval notation, this is just \[ \forall t \in [0,1]\cap [2, \infty) \quad v'(t) > 0 \]

OpenStudy (anonymous):

@kizi239 Think you got it?